If $51.30 interest is earned in two years on a deposit of $95, then what is the simple annual interest rate?

1
Expert's answer
2022-01-27T16:12:04-0500

If a principal of P dollars is deposited for T years at a simple interest rate of R, then the interest, I ,will be


I = P*R*T


In our case;


I = $51.30

P = $95

T= 2 years

R = ?


From I = P*R*T we get


R = I/(P*T)


R = 51.30/95*2


R = 0.27 or 0.27 *100 = 27% interest